OpenStudy (anonymous):

Theory number. n . (2^n) is perfect square number with n is elements of set {1 , 2, 3, 4, ... , 2015}. How many n which satisfy this condition.

OpenStudy (anonymous):

looks n is even number , right ?

Miracrown (miracrown):

Yes, n would have to be an even number. That is necessary, but not sufficient.

Miracrown (miracrown):

I think n also have to be a perfect square itself.

Miracrown (miracrown):

The odd factors definitely have to be squared.

Miracrown (miracrown):

Let me think about the factors of 2 in n.

OpenStudy (anonymous):

n = 2^k with k = 2,4,6, ...

Miracrown (miracrown):

Having zero powers of 2 does not work. Having one power of two does not work. Having two powers of two does work.

Miracrown (miracrown):

No, n = p^2 * 2^k with k = 2, 4, 6... where p is an odd number.

Miracrown (miracrown):

So let's take k = 2 first, 2^2 = 4 2015/4 = 503.75 Then we want how many odd perfect squares less than 503.75 sqrt(503.75) = 22.44 So that looks like we will have odd squares for 1 through 21 That's 11, yes?

OpenStudy (anonymous):

looks yours is same with mine n . 2^n = p^2 for n an even number so 2^n is already be perfect, therefore n also must be perfect square. because perfect square times perfect square is perfect square.